OpenStudy (konradzuse):

evaluate each integral (csc^2 * t) / (cot t) dx

OpenStudy (anonymous):

write it by usgin equation plz so we can read : )

OpenStudy (anonymous):

it can be written 1/sin(x)cos(x) 2csc(2x) now take out integral ln(csc2x-cot2x)+c

OpenStudy (anonymous):

\[\frac{1}{sin^2t}*\frac{sin t}{cost}\implies \frac{1}{sint *cost}\]\[t = sinx\]\[dx=\frac{dt}{cosx}\] \[\int\frac{1}{t(1-t^2)}dt\]

OpenStudy (anonymous):

Now Integrate it using partial fraction

OpenStudy (anonymous):

I used this idetity sin(2x)=2sinxcosx

OpenStudy (konradzuse):

okay so for some reason this site is very slow and myequation was deleted :(

OpenStudy (anonymous):

since d/dt(cot(t))=-csc^2t \[\int\limits_{}^{}-1/\cot(t)d(\cot(t))=\int\limits_{}^{}\csc^2(t)/\cot(d)d\] so since the derivative of cot(t) is sitting in the integral you can let "u" be cot(t) INT 1/u=ln(u)=-ln(cot(t)) you add a (-) sign because the d/dt(cot(t))=-csc^2t answer=-ln(cot(t)) +C

OpenStudy (anonymous):

you meant dt at the top right not dx?

OpenStudy (konradzuse):

dt yes I messed up.

OpenStudy (anonymous):

There’s more than one way to skin a cat : )

OpenStudy (anonymous):

yes, do you understand what i mean when i say \[\int\limits_{}^{}\csc^2(t)/\cot(t)dt=\int\limits_{}^{}-1/\cot(t)d(\cot(t))dt\]